OpenStudy (anonymous):

draw a sketch of an angle in standard position having the equation 3x-4y=0, x is less than or equal to 0, as its terminal side. indicate the angle of least positive measure theta, and give the values of sin theta, cos theta, tan theta, cot theta, sec theta,and csc theta

ganeshie8 (ganeshie8):

rewrite the given equation of line as : \[4y=3x\\y=\frac{3}{4}x\]

ganeshie8 (ganeshie8):

do you know how to sketch that line ?

OpenStudy (anonymous):

no and how did you get \[\frac{ 3 }{ 4 }x\]

ganeshie8 (ganeshie8):

divided 4 both sides

OpenStudy (anonymous):

okay! got that part!

ganeshie8 (ganeshie8):

good, can you sketch that line ?

ganeshie8 (ganeshie8):

the line is of form \(y=mx\) so it passes through origin

OpenStudy (anonymous):

no :( and okay

ganeshie8 (ganeshie8):

y = 3/4x that means the line rises 3 units for a run of 4 units 

ganeshie8 (ganeshie8):

they want you consider the line only for x < 0 so, it would be just this :